40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11269     Entry time: Sun May 3 19:40:51 2015
Author: rana 
Type: Update 
Category: ASC 
Subject: Sunday maintenance: alignment, OL center, seismo, temp sensors 

X arm was far out in yaw, so I reran the ASS for Y and then X. Ran OK; the offload from ASS outputs to SUS bias is still pretty violent - needs smoother ramping.

After this I recentered the ITMX OL- it was off by 50 microradians in pitch. Just like the BS/PRM OLs, this one has a few badly assembled & flimsly mounts. Steve, please prepare for replacing the ITMX OL mirror mounts with the proper base/post/Polaris combo. I think we need ~3 of them. Pit/yaw loop measurements attached.

Based on the PEM-SEIS summary page, it looked like GUR1 was oscillating (and thereby saturating and suppressing the Z channel). So I power cycled both Guralps by turning off the interface box for ~30 seconds and the powering back on. Still not fixed; looks like the oscillations at 110 and 520 Hz have moved but GUR2_X/Y are suppressed above 1 Hz, and GUR1_Z is suppressed below 1 Hz. We need Jenne or Zach to come and use the Gur Paddle on these things to make them OK.

From the SUS-WatchDog summary page, it looked like the PRM tripped during the little 3.8 EQ at 4AM, so I un-tripped it.

Caryn's temperature sensors look like they're still plugged in. Does anyone know where they're connected?

Attachment 1: itmx_ol_loops_150503.png  68 kB  Uploaded Sun May 3 21:47:08 2015  | Hide | Hide all
itmx_ol_loops_150503.png
Attachment 2: Gur_150503.png  198 kB  Uploaded Sun May 3 22:07:45 2015  | Hide | Hide all
Gur_150503.png
ELOG V3.1.3-